# Resolving Quantum Criticality in the Honeycomb Hubbard Model

### v1

@@
The interplay between Dirac fermions and electronic correlations on the honeycomb lattice hosts a fundamental quantum phase transition from a semimetal to a Mott insulator, governed by the Gross-Neveu-Heisenberg (GNH) universality class. Despite its importance, consensus on the precise critical exponents remains elusive due to severe finite-size effects in numerical simulations and the lack of conformal bootstrap benchmarks. Here we try to resolve this long-standing controversy by performing projector determinant quantum Monte Carlo (QMC) simulations on lattices of unprecedented size, reaching 10,368 sites. By developing a novel projected submatrix update algorithm, we achieve a significant algorithmic speedup that enables us to access the thermodynamic limit with high precision. We observe that the fermion anomalous dimension and the correlation length exponent converge rapidly, while the boson anomalous dimension exhibits a systematic size dependence that we resolve via linear extrapolation. To validate our analysis, we perform parallel large-scale simulations of the spinless $t$-$V$ model on the honeycomb lattice, which belongs to the Gross-Neveu-Ising class. Our results for the $t$-$V$ model, including the first QMC determination of the fermion anomalous dimension, show agreement with conformal bootstrap predictions, thereby corroborating the robustness of our methodology. Our work provides state-of-the-art critical exponents for the honeycomb Hubbard model and establishes a systematic finite-size scaling workflow applicable to a broad class of strongly correlated quantum systems, paving the way for resolving other challenging fermionic quantum critical phenomena.
@@ I Introduction
Understanding how quantum materials transition between distinct phases of matter is a central challenge in condensed matter physics. When massless Dirac fermions–the relativistic quasiparticles underlying graphene’s electronic properties and emerging in diverse quantum materials from $d$-wave superconductors to surface of topological insulators [1, 2, 3]–interact strongly, they can undergo quantum phase transitions that lie beyond conventional theoretical paradigms. A prototypical example is the interaction-driven semimetal-to-Mott-insulator transition in the honeycomb Hubbard model [4, 5, 6, 7, 8, 9], where increasing electron-electron repulsion spontaneously generates mass and magnetic order. This transition belongs to the Gross-Neveu-Heisenberg (GNH) universality class [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and connects fundamental questions in quantum materials to high-energy physics concepts such as chiral symmetry breaking and dynamical mass generation.
@@
Despite decades of effort, consensus on the precise critical exponents governing this transition remains elusive, as illustrated in Fig. 1. Analytical approaches–including $\epsilon$ expansions, large-$N$ expansions, and functional renormalization group (FRG) calculations–yield estimates that are only partially compatible, with a particularly pronounced spread in the fermion anomalous dimension $\eta_{\psi}$ [10, 11, 12, 13]. Quantum Monte Carlo (QMC) results are even less uniform: for the honeycomb Hubbard model, most studies report a boson anomalous dimension $\eta_{\phi}\lesssim 0.65$ [14, 15, 16, 17, 18]–approximately 40% smaller than many analytical predictions [10, 11, 12, 13]–while different lattice realizations in the same universality class yield noticeably larger values. In contrast, for the $N=4$ Gross-Neveu-Ising (GNI) universality class [26, 27, 28, 29, 30, 31, 15, 32, 33, 34, 35, 36], large-scale QMC simulations [36] have produced critical exponents consistent with conformal bootstrap predictions [31], highlighting the achievable precision when finite-size effects are properly controlled.
@@
Figure 1: Summary of critical exponents $\nu$, $\eta_{\phi}$, and $\eta_{\psi}$ reported in the literature and obtained in this work. a, $N=8$ Gross-Neveu-Heisenberg universality class [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. b, $N=4$ Gross-Neveu-Ising universality class [26, 27, 28, 29, 30, 31, 15, 32, 33, 34, 35, 36]. The gray horizontal dashed lines divide the plot into three blocks: analytical results (top), QMC results for the same model as in this work (middle), and QMC results for other models in the same universality class (bottom). CB denotes conformal bootstrap [31]. See Supplemental Material Tables S5 and S6 for the numerical values and methodological details.
@@
The root cause of the GNH controversy lies in severe finite-size effects that render extracted exponents highly sensitive to the choice of system sizes and fitting procedures. The small leading-correction-to-scaling exponent $\omega\approx 0.3$ arises from irrelevant operators associated with competing interaction channels [11] and non-universal analytic contributions grounded in the large $\eta_{\phi}$ [19], which also induce large statistical fluctuations in susceptibilities sampled from QMC [22]. Alternative simulation setups have been proposed to mitigate these effects, including extended-hopping models that enlarge the linear dispersion region [37, 38, 39, 20] and elective-momentum ultrasize QMC methods [40]. Remarkably, SLAC-fermion formulations produce high-precision exponents and suggest that honeycomb estimates drift toward SLAC values once the smallest system sizes are excluded [20], though a noticeable discrepancy with analytical predictions persists. However, the honeycomb Hubbard model remains the minimal model for the GNH transition with direct experimental relevance to graphene and related two-dimensional materials, making it essential to resolve the controversy within this original setting. While hybrid Monte Carlo (HMC) simulations [16] have achieved lattices with $L>100$, ergodicity issues restrict reliable results to high temperatures, preventing direct access to ground-state properties. Their extracted correlation length exponent $\nu$ is consistent with theoretical predictions, but $\eta_{\phi}$ still deviates significantly from analytical estimates, and $\eta_{\psi}$ remains undetermined [16].
@@
In this work, we revisit this long-standing problem with the numerically exact projector determinant quantum Monte Carlo (PQMC) [41, 42, 43] method on lattices up to $72\times 72\times 2$ sites—twice the linear dimension of previous PQMC study [18]. This unprecedented scale is enabled by a new submatrix update algorithm that optimizes cache utilization in CPU operations [44, 45, 46], achieving state-of-the-art performance without introducing any additional approximations. Through a straightforward yet systematic finite-size scaling (FSS) analysis, we unveil qualitatively distinct convergence patterns for the three critical exponents and extract their definitive values accordingly. Parallel simulations of the spinless $t$-$V$ model belonging to $N=4$ GNI universality class validate our approach and demonstrate that scaling up to such large sizes reveals the characteristic finite-size behaviors essential for resolving fermionic quantum criticality.
@@ II.1 Semimetal-Mott-insulator transition in the honeycomb Hubbard model
We study the half-filled Hubbard model on the honeycomb lattice. The Hamiltonian is given by $\hat{H}=-t\sum_{\langle i,j\rangle,\sigma}(c_{i\sigma}^{\dagger}c_{j\sigma}+\mathrm{H.c.})+U\sum_{i}\left(\hat{n}_{i\uparrow}-\frac{1}{2}\right)\left(\hat{n}_{i\downarrow}-\frac{1}{2}\right)$, where $t$ is the nearest-neighbor hopping amplitude and $U$ is the Hubbard interaction strength. We impose periodic boundary conditions on an $L\times L$ honeycomb lattice with two sublattices, yielding $N_{\mathrm{s}}=2L^{2}$ lattice sites. At half filling, the low-energy excitations are spin-1/2 interacting Dirac fermions at the Dirac points $\mathrm{K}$ and $\mathrm{K^{\prime}}$. Increasing $U$ drives a continuous Mott transition belonging to the $N=8$ GNH universality class, from a semimetal to an antiferromagnetic (AFM) insulator with spontaneous spin-rotation symmetry breaking [4, 6, 7, 8, 9, 14, 15, 16, 17, 18, 19, 20], as illustrated in Fig. 2a. We employ the sign-problem-free PQMC [41, 42, 43] to obtain the ground-state properties of this model, with the electron number fixed to $N_{\mathrm{s}}$ to enforce half filling. The formulation, optimization, implementation, and parameter choices of the algorithm are provided in Methods and Supplemental Material [47].
@@
For finite systems, the crossing point between a pair of system sizes exhibits a drift due to scaling corrections [49, 19]. As shown in Fig. 2b, only for $L\gtrsim 36$ do the $R_{\mathrm{AFM}}$-$U$ curves approximately cross at a common point, indicating strong finite-size effects in the honeycomb Hubbard model. The final crossing point ($\sim 3.74$) is consistent with the transition scale estimated from the decay of the real-space spin-spin correlation function at maximum distance (see Supplemental Material [47] Section S6). To quantify this drift, we denote the crossing point between sizes $L$ and $2L$ as $U^{\times}(L,2L)$ (horizontal bars in Fig. 2b) and fit $U^{\times}(L,2L)=U_{c}^{\times}+a(2L)^{-b}$. The inset of Fig. 2b shows that this power-law extrapolation yields $U_{c}^{\times}=3.693(7)$ and $b_{\mathrm{fit}}=1.26(5)$. For an RG-invariant quantity, the drift exponent is expected to scale as $b=1/\nu+\min(\omega,2-z-\eta_{\phi})$ [49], where $\omega$ is the correction-to-scaling exponent from irrelevant operators and $z=1$ for the Lorentz-invariant Gross-Neveu criticality. Using the data-collapse estimates $\nu=1.11(9)$ and $\eta_{\phi}=0.715(15)$ obtained later with the data from the largest sizes, and noting that one-loop $\epsilon$-expansion studies suggest $\omega\sim 0.3$ for the $N=8$ GNH fixed point [11], we obtain $b_{\mathrm{collapse}}=1.19(5)$, consistent with $b_{\mathrm{fit}}$ within error bars.
@@ II.2 Gross-Neveu-Heisenberg criticality
Figure 3: Sliding-window data-collapse analysis for the Hubbard model. Columns show three representative fit windows with $L_{\mathrm{max}}=36$, 54, and 72 (see Supplemental Material [47] Section S6 for all windows). The $L$-axis above each panel indicates the sizes included in the fit, with greyed-out sizes excluded. Rows: $R_{\mathrm{AFM}}^{(1,0)}$ collapse yielding $(U_{c},\nu)$ (top), $m^{2}_{\mathrm{AFM}}$ collapse yielding $\eta_{\phi}$ (middle), and $\langle|G_{AB}|\rangle$ collapse yielding $\eta_{\psi}$ (bottom).
@@
We provide an argument for this ansatz, including the rationale for taking the modulus, in Supplemental Material [47] Section S5.
@@
Given the extended range of accessible system sizes, selecting an appropriate fitting window is not obvious a priori. A common strategy in previous QMC studies is to successively discard the smallest sizes until $\chi^{2}_{\mathrm{red}}$ (see Methods) becomes acceptable [19, 23, 48, 20]. We refer to this practice as a shrinking-window analysis: the window width (i.e. the number of consecutive system sizes included in the fit) decreases while the largest size $L_{\max}$ is fixed. Notably, Ref. [20] reported a monotonic drift of $\eta_{\phi}$ with increasing the smallest size $L_{\min}$, trending toward the SLAC-fermion estimate. To learn this window dependence systematically, we scan over both upper and lower cutoffs $L_{\min}$ and $L_{\max}$, finding that overly narrow windows usually yield larger error bars (since data are sparser and thus impose fewer constraints on the fit), whereas overly wide windows tend to increase $\chi^{2}_{\mathrm{red}}$ (see also the tables in Refs. [19, 48]). The complete scan over fitting windows is shown in the heatmaps in Supplemental Material [47] Section S9.
Based on the above observations, we adopt a sliding-window data-collapse analysis: we fix the window width at six consecutive sizes and slide it across all accessible $L$ (choosing a window width of five yields similar results). Within each window, we fit the scaling forms in Eqs. (4)–(6) and extract $(U_{c},\nu,\eta_{\phi},\eta_{\psi})$ following Methods, labeling each window by its largest size $L_{\mathrm{max}}$. Figure 3 illustrates this procedure for the Hubbard model: as the window slides from left to right toward larger $L$, the fitted $U_{c}$, $\eta_{\phi}$, and $\eta_{\psi}$ drift monotonically, whereas $\nu$ appears to oscillate. For the $L_{\mathrm{max}}=72$ window (rightmost column), we obtain $U_{c}=3.7422(34)$, $\nu=1.11(5)$, $\eta_{\phi}=0.715(15)$, and $\eta_{\psi}=0.1881(33)$. We also perform the same FSS analysis for the spinless $t$-$V$ model on the honeycomb lattice, whose critical point belongs to the $N=4$ GNI universality class, with system sizes up to $L=48$ (see Supplemental Material [47] Section S7 for details).
@@ III Discussion
Placed in the broader context of prior studies, our results are compared with representative analytical calculations and large-scale QMC simulations in Fig. 1. For the $N=4$ GNI transition realized in the spinless $t$-$V$ model, our $\nu$ and $\eta_{\phi}$ agree with both conformal bootstrap (CB) predictions [31] and fermion-bag QMC simulations [36]. Most notably, we provide the first QMC estimate of $\eta_{\psi}$ in this model, which is consistent with the CB prediction within uncertainties. For the $N=8$ GNH transition in the honeycomb Hubbard model, the wide confidence interval of our $\nu$ encompasses most prior QMC and analytical values. Our $\eta_{\phi}$ moves closer to analytical predictions and aligns with most other lattice realizations in the same universality class [25, 24, 23, 22, 20], and our $\eta_{\psi}\approx 0.19$ remains consistent with earlier QMC values near $0.2$, though a gap to analytical predictions persists for both anomalous dimensions.
The sliding-window analysis in Fig. 4 offers internal validation of our final estimates, by tracking how the fitted parameters evolve as system size increases. When we restrict the Hubbard data to $L\leq 36$, from Figs. 3a, d and g, we obtain $U_{c}=3.8187(30)$, $\nu=1.058(27)$, $\eta_{\phi}=0.628(7)$, and $\eta_{\psi}=0.1948(15)$, values closer to earlier PQMC studies on the same model [18, 20]. As the sliding window moves to larger sizes, $\eta_{\phi}$ increases monotonically, and $\eta_{\psi}$ decreases slightly, all shifting from earlier honeycomb Hubbard estimates toward analytical predictions. This systematic drift demonstrates that finite-size effects are the primary source of discrepancies between earlier honeycomb Hubbard results and other estimates. More broadly, beyond serving as an internal consistency check in the present work, the sliding-window analysis offers a transparent way to assess the stability of FSS estimates as system size grows. Compared with the conventional shrinking-window strategy, where $L_{\max}$ is fixed and $L_{\min}$ is thereby bounded from above, the sliding-window shifts the entire fit interval toward larger sizes and thus provides a more natural extrapolation to the TDL.
@@
Looking ahead, the sliding-window protocol developed here can prove broadly useful for extracting critical exponents in other quantum phase transitions, especially when finite-size effects are severe. With both the computational power to reach larger sizes and a robust protocol to extract asymptotic exponents, the path toward resolving other outstanding problems in strongly correlated quantum systems–particularly those involving competing orders, emergent gauge structures, or unconventional quantum criticality–is now significantly clearer.

### v2

@@
Quantum phase transitions driven by electronic correlations are central to understanding the physics of graphene and related two-dimensional materials. A paradigmatic example is the semimetal-to-Mott-insulator transition on the honeycomb lattice, governed by the Gross-Neveu-Heisenberg universality class, yet consensus on its critical exponents has remained elusive for over a decade due to severe finite-size effects and the absence of rigorous conformal bootstrap benchmarks. Here we try to resolve this long-standing controversy by performing projector determinant quantum Monte Carlo simulations on lattices of unprecedented size, reaching 10,368 sites. By developing a novel projected submatrix update algorithm, we achieve a significant algorithmic speedup that enables us to access the thermodynamic limit with high precision. We observe that the fermion anomalous dimension and the correlation length exponent converge rapidly, while the boson anomalous dimension exhibits a systematic size dependence that we resolve via linear extrapolation. To validate our analysis, we perform parallel large-scale simulations of the spinless $t$-$V$ model on the honeycomb lattice, which belongs to the Gross-Neveu-Ising class. Our results for the $t$-$V$ model show agreement with conformal bootstrap predictions, thereby corroborating the robustness of our methodology. Our work provides state-of-the-art critical exponents for the honeycomb Hubbard model and establishes a systematic finite-size scaling workflow applicable to a broad class of strongly correlated quantum systems, paving the way for resolving other challenging fermionic quantum critical phenomena.
@@ I Introduction
Understanding how quantum materials transition between distinct phases of matter is a central challenge in condensed matter physics. When massless Dirac fermions–the relativistic quasiparticles underlying graphene’s electronic properties and emerging in diverse quantum materials from $d$-wave superconductors to surface of topological insulators [1, 2]–interact strongly, they can undergo quantum phase transitions that lie beyond conventional theoretical paradigms. A prototypical example is the interaction-driven semimetal-to-Mott-insulator transition in the honeycomb Hubbard model [3, 4, 5, 6, 7, 8], where increasing electron-electron repulsion spontaneously generates mass and magnetic order. This relativistic Mott transition has recently been observed experimentally in an artificial graphene realized by twisted double bilayer WSe2 [9, 10, 11], in which the twist angle tunes the interaction strength. The associated quantum critical point belongs to the Gross-Neveu-Heisenberg (GNH) universality class [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] and connects fundamental questions in quantum materials to high-energy physics concepts such as chiral symmetry breaking and dynamical mass generation.
@@
Despite decades of effort, consensus on the precise critical exponents governing this transition remains elusive, as illustrated in Fig. 1. Analytical approaches–including $\epsilon$ expansions, large-$N$ expansions, and functional renormalization group (FRG) calculations–yield estimates that are only partially compatible, with a particularly pronounced spread in the fermion anomalous dimension $\eta_{\psi}$ [12, 13, 14, 16]. Quantum Monte Carlo (QMC) results are even less uniform: for the honeycomb Hubbard model, most studies report a boson anomalous dimension $\eta_{\phi}\lesssim 0.65$ [17, 18, 19, 20, 21]–approximately 40% smaller than many analytical predictions [12, 13, 14, 16]–while different lattice realizations in the same universality class yield noticeably larger values. In contrast, for the $N=4$ Gross-Neveu-Ising (GNI) universality class [29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38, 39, 40, 41], large-scale QMC simulations [41] have produced critical exponents consistent with conformal bootstrap predictions [35], highlighting the achievable precision when finite-size effects are properly controlled.
@@
Figure 1: Summary of critical exponents $\nu$, $\eta_{\phi}$, and $\eta_{\psi}$ reported in the literature and obtained in this work. a, $N=8$ Gross-Neveu-Heisenberg universality class [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. b, $N=4$ Gross-Neveu-Ising universality class [29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38, 39, 40, 41]. The gray horizontal dashed lines divide the plot into three blocks: analytical results (top), QMC results for the same model as in this work (middle), and QMC results for other models in the same universality class (bottom). CB denotes conformal bootstrap [35]. See Supplementary Tables 5 and 6 for the numerical values and methodological details.
@@
The root cause of the GNH controversy lies in severe finite-size effects that render extracted exponents highly sensitive to the choice of system sizes and fitting procedures. The small leading-correction-to-scaling exponent $\omega\approx 0.3$ arises from irrelevant operators associated with competing interaction channels [13] and non-universal analytic contributions grounded in the large $\eta_{\phi}$ [22], which also induce large statistical fluctuations in susceptibilities sampled from QMC [25]. Alternative simulation setups have been proposed to mitigate these effects, including extended-hopping models that enlarge the linear dispersion region [42, 43, 44, 23] and elective-momentum ultrasize QMC methods [45]. Remarkably, SLAC-fermion formulations produce high-precision exponents and suggest that honeycomb estimates drift toward SLAC values once the smallest system sizes are excluded [23], though a noticeable discrepancy with analytical predictions persists. However, the honeycomb Hubbard model remains the minimal model for the GNH transition with direct experimental relevance to graphene and related two-dimensional materials, making it essential to resolve the controversy within this original setting. While hybrid Monte Carlo (HMC) simulations [19] have achieved lattices with $L>100$, ergodicity issues restrict reliable results to high temperatures, preventing direct access to ground-state properties. Their extracted correlation length exponent $\nu$ falls within the range spanned by various analytical estimates, yet the boson anomalous dimension $\eta_{\phi}$ remains notably smaller than most field-theoretic predictions, and $\eta_{\psi}$ was not determined [19].
@@
In this work, we revisit this long-standing problem with the numerically exact projector determinant quantum Monte Carlo (PQMC) [46, 47, 48] method on lattices up to $72\times 72\times 2$ sites—twice the linear dimension of previous PQMC study [21]. This unprecedented scale is enabled by a new submatrix update algorithm that optimizes cache utilization in CPU operations, achieving state-of-the-art performance without introducing any additional approximations. Through a straightforward yet systematic finite-size scaling (FSS) analysis, we unveil qualitatively distinct convergence patterns for the three critical exponents and extract their definitive values accordingly. Parallel simulations of the spinless $t$-$V$ model belonging to $N=4$ GNI universality class validate our approach and demonstrate that scaling up to such large sizes reveals the characteristic finite-size behaviors essential for resolving fermionic quantum criticality.
@@ II.1 Semimetal-Mott-insulator transition in the honeycomb Hubbard model
We study the half-filled Hubbard model on the honeycomb lattice. The Hamiltonian is given by $\hat{H}=-t\sum_{\langle i,j\rangle,\sigma}(c_{i\sigma}^{\dagger}c_{j\sigma}+\mathrm{H.c.})+U\sum_{i}\left(\hat{n}_{i\uparrow}-\frac{1}{2}\right)\left(\hat{n}_{i\downarrow}-\frac{1}{2}\right)$, where $t$ is the nearest-neighbor hopping amplitude and $U$ is the Hubbard interaction strength. We impose periodic boundary conditions on an $L\times L$ honeycomb lattice with two sublattices, yielding $N_{\mathrm{s}}=2L^{2}$ lattice sites. At half filling, the low-energy excitations are spin-1/2 interacting Dirac fermions at the Dirac points $\mathrm{K}$ and $\mathrm{K^{\prime}}$. Increasing $U$ drives a continuous Mott transition belonging to the $N=8$ GNH universality class, from a semimetal to an antiferromagnetic (AFM) insulator with spontaneous spin-rotation symmetry breaking [3, 5, 6, 7, 8, 17, 18, 19, 20, 21, 22, 23], as illustrated in Fig. 2a. We employ the sign-problem-free PQMC [46, 47, 48] to obtain the ground-state properties of this model, with the electron number fixed to $N_{\mathrm{s}}$ to enforce half filling. The formulation, optimization, implementation, and parameter choices of the algorithm are provided in Methods and Supplementary Information.
@@
For finite systems, the crossing point between a pair of system sizes exhibits a drift due to scaling corrections [50, 22]. As shown in Fig. 2b, only for $L\gtrsim 36$ do the $R_{\mathrm{AFM}}$-$U$ curves approximately cross at a common point, indicating strong finite-size effects in the honeycomb Hubbard model. The final crossing point ($\sim 3.74$) is consistent with the transition scale estimated from the decay of the real-space spin-spin correlation function at maximum distance (see Supplementary Note 6). To quantify this drift, we denote the crossing point between sizes $L$ and $2L$ as $U^{\times}(L,2L)$ (horizontal bars in Fig. 2b) and fit $U^{\times}(L,2L)=U_{c}^{\times}+a(2L)^{-b}$. The inset of Fig. 2b shows that this power-law extrapolation yields $U_{c}^{\times}=3.693(7)$ and $b_{\mathrm{fit}}=1.26(5)$. For an RG-invariant quantity, the drift exponent is expected to scale as $b=1/\nu+\min(\omega,2-z-\eta_{\phi})$ [50], where $\omega$ is the correction-to-scaling exponent from irrelevant operators and $z=1$ for the Lorentz-invariant Gross-Neveu criticality. Using the data-collapse estimates $\nu=1.11(9)$ and $\eta_{\phi}=0.715(15)$ obtained later with the data from the largest sizes, and noting that one-loop $\epsilon$-expansion studies suggest $\omega\sim 0.3$ for the $N=8$ GNH fixed point [13], we obtain $b_{\mathrm{collapse}}=1.19(5)$, consistent with $b_{\mathrm{fit}}$ within error bars.
@@ II.2 Gross-Neveu-Heisenberg criticality
Figure 3: Sliding-window data-collapse analysis for the Hubbard model. Columns show three representative fit windows with $L_{\mathrm{max}}=36$, 54, and 72 (see Supplementary Note 6 for all windows). The $L$-axis above each panel indicates the sizes included in the fit, with greyed-out sizes excluded. Rows: $R_{\mathrm{AFM}}^{(1,0)}$ collapse yielding $(U_{c},\nu)$ (top), $m^{2}_{\mathrm{AFM}}$ collapse yielding $\eta_{\phi}$ (middle), and $\langle|G_{AB}|\rangle$ collapse yielding $\eta_{\psi}$ (bottom).
@@
We provide an argument for this ansatz, including the rationale for taking the modulus, in Supplementary Note 5.
@@
Given the extended range of accessible system sizes, selecting an appropriate fitting window is not obvious a priori. A common strategy in previous QMC studies is to successively discard the smallest sizes until $\chi^{2}_{\mathrm{red}}$ (see Methods) becomes acceptable [22, 26, 49, 23]. We refer to this practice as a shrinking-window analysis: the window width (i.e. the number of consecutive system sizes included in the fit) decreases while the largest size $L_{\max}$ is fixed. Notably, Ref. [23] reported a monotonic drift of $\eta_{\phi}$ with increasing the smallest size $L_{\min}$, trending toward the SLAC-fermion estimate. To learn this window dependence systematically, we scan over both upper and lower cutoffs $L_{\min}$ and $L_{\max}$, finding that overly narrow windows usually yield larger error bars (since data are sparser and thus impose fewer constraints on the fit), whereas overly wide windows tend to increase $\chi^{2}_{\mathrm{red}}$ (see also the tables in Refs. [22, 49]). The complete scan over fitting windows is shown in the heatmaps in Supplementary Note 8.
Based on the above observations, we adopt a sliding-window data-collapse analysis: we fix the window width at six consecutive sizes and slide it across all accessible $L$ (choosing a window width of five yields similar results). Within each window, we fit the scaling forms in Eqs. (4)–(6) and extract $(U_{c},\nu,\eta_{\phi},\eta_{\psi})$ following Methods, labeling each window by its largest size $L_{\mathrm{max}}$. Figure 3 illustrates this procedure for the Hubbard model: as the window slides from left to right toward larger $L$, the fitted $U_{c}$, $\eta_{\phi}$, and $\eta_{\psi}$ drift monotonically, whereas $\nu$ appears to oscillate. For the $L_{\mathrm{max}}=72$ window (rightmost column), we obtain $U_{c}=3.7422(34)$, $\nu=1.11(5)$, $\eta_{\phi}=0.715(15)$, and $\eta_{\psi}=0.1881(33)$. We also perform the same FSS analysis for the spinless $t$-$V$ model on the honeycomb lattice, whose critical point belongs to the $N=4$ GNI universality class, with system sizes up to $L=48$ (see Supplementary Note 7 for details).
@@ III Discussion
Placed in the broader context of prior studies, our results are compared with representative analytical calculations and large-scale QMC simulations in Fig. 1. For the $N=4$ GNI transition realized in the spinless $t$-$V$ model, our $\nu$ and $\eta_{\phi}$ agree with both conformal bootstrap (CB) predictions [35] and fermion-bag QMC simulations [41]. Most notably, our estimate of $\eta_{\psi}$ in this model presents the first QMC determination that is consistent with the CB prediction within uncertainties. For the $N=8$ GNH transition in the honeycomb Hubbard model, the wide confidence interval of our $\nu$ encompasses most prior QMC and analytical values. Our $\eta_{\phi}$ moves closer to analytical predictions and aligns with most other lattice realizations in the same universality class [28, 27, 26, 25, 23], and our $\eta_{\psi}\approx 0.19$ remains consistent with earlier QMC values near $0.2$, though some tension with analytical estimates persists for both anomalous dimensions.
The sliding-window analysis in Fig. 4 offers internal validation of our final estimates, by tracking how the fitted parameters evolve as system size increases. When we restrict the Hubbard data to $L\leq 36$, from Figs. 3a, d and g, we obtain $U_{c}=3.8187(30)$, $\nu=1.058(27)$, $\eta_{\phi}=0.628(7)$, and $\eta_{\psi}=0.1948(15)$, values closer to earlier PQMC studies on the same model [21, 23]. As the sliding window moves to larger sizes, $\eta_{\phi}$ increases monotonically, and $\eta_{\psi}$ decreases slightly, all shifting from earlier honeycomb Hubbard estimates toward analytical predictions. This systematic drift demonstrates that finite-size effects are the primary source of discrepancies between earlier honeycomb Hubbard results and other estimates. More broadly, beyond serving as an internal consistency check, the sliding-window analysis offers a transparent way to assess the stability of FSS estimates as system size grows. Compared with the conventional shrinking-window strategy, where $L_{\max}$ is fixed and $L_{\min}$ is thereby bounded from above, the sliding-window shifts the entire fit interval toward larger sizes and thus provides a more natural extrapolation to the TDL.
@@
Looking ahead, the sliding-window protocol developed in this work can prove broadly useful for extracting critical exponents in other quantum phase transitions, especially when finite-size effects are severe. With both the computational power to reach larger sizes and a robust protocol to extract asymptotic exponents, the path toward resolving other outstanding problems in strongly correlated quantum systems–particularly those involving competing orders, emergent gauge structures, or unconventional quantum criticality–is now significantly clearer.
Further progress in the GNH criticality will also benefit from independent benchmarks from experiments and from CB. In particular, measurements of critical exponents in artificial graphene realized by moiré materials [9] may become possible in the future. On the theory side, extending CB constraints available for the $O(N)$ Gross-Neveu-Yukawa model [35] to the GNH fixed point would offer a complementary nonperturbative reference.